OpenStudy (anonymous):

Multiply x^2-49/x^2*x^2-7x/x^2+3x-70

OpenStudy (anonymous):

is that how it is written?

OpenStudy (anonymous):

I might be missing somthing let me check lol

OpenStudy (anonymous):

(x^2-49)/(x^2)*(x^2-7x)/(x^2+3x-70)

OpenStudy (anonymous):

\[\LARGE \frac{x^2-49}{x^2}\cdot \frac{x^2-7x}{x^2+3x-70 }\] \[\LARGE \frac{x^2-7^2}{x^2}\cdot \frac{x(x-7)}{x^2+3x-70 }\] \[\LARGE \frac{(x-7)(x+7 )}{x^{\cancel{2}}}\cdot \frac{\cancel{x} (x-7)}{x^2+3x-70 }\] .... try factorizing the denominator of the second fraction...

OpenStudy (anonymous):

\[\LARGE \frac{(x-7)(x+7 )}{x}\cdot \frac{ (x-7)}{(x+10)(x-7 ) }\] \[\LARGE \frac{(x-7)(x+7 )}{x}\cdot \frac{ \cancel{(x-7)}}{(x+10)\cancel{(x-7 )} }\]

OpenStudy (anonymous):

now multiply them...

OpenStudy (anonymous):

since your question says... Multiply x^2-49/x^2*x^2-7x/x^2+3x-70